OpenStudy (anonymous):

whats 6/8 divided by 2/8

OpenStudy (anonymous):

6/8 * 8/2

OpenStudy (anonymous):

do i cross multiply

OpenStudy (anonymous):

yep and then simplify

OpenStudy (anonymous):

ow do you simplify

OpenStudy (anonymous):

is it 12/16

OpenStudy (anonymous):

(6*8)/*(8*2)

OpenStudy (anonymous):

i dont know how to simplify

OpenStudy (anonymous):

divide both numerator and denomitaor with 4 since 4 is the common number

OpenStudy (anonymous):

so 12/16 becomes 3/4

OpenStudy (asnaseer):

the rule to follow here is:\[\frac{a}{b}\div\frac{c}{d}=\frac{a}{b}\times\frac{d}{c}\]

OpenStudy (asnaseer):

so, in your case, we have:\[\frac{6}{8}\div\frac{2}{8}=\frac{6}{\cancel{8}}\times\frac{\cancel{8}}{2}\]

OpenStudy (asnaseer):

@haileyshay can you work it out now?

Parth (parthkohli):

I agree, but we can also do something like \[ {a \over c} \div {b \over d} = {a \div b \over c \div d} \Longrightarrow {6 \div 2 \over 8 \div 8}\] :)

OpenStudy (anonymous):

(6*8)/*(8*2) = 48/16 = 3

Parth (parthkohli):

That gives us the computation directly.

OpenStudy (anonymous):

wats 3/5 divded by 3/8

OpenStudy (asnaseer):

use the same process

OpenStudy (asnaseer):

why don't you try it step-by-step

OpenStudy (anonymous):

i did it said its wrong

OpenStudy (asnaseer):

please show each step you took - then I'll try and help you spot where you may have made a mistake.

OpenStudy (anonymous):

3x8

OpenStudy (asnaseer):

that was your first step?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

try using this form:\[\frac{a}{b}\div\frac{c}{d}=\frac{a}{b}\times\frac{d}{c}\]now plug-in your values and see what you get.

OpenStudy (anonymous):

so 3 divided by 3

OpenStudy (asnaseer):

can you please write out your step fully using the formula I gave you above?

OpenStudy (asnaseer):

your question is: 3/5 divided by 3/8 so in this case: a=3 b=5 c=3 d=8 plug those into the formula I gave you.

OpenStudy (anonymous):

i dont know how to make it look like a fraction on the computer

OpenStudy (asnaseer):

try using the Draw button and draw it out if that is easier.

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

e.g.: